OpenStudy (anonymous):

how to integrate (t^3-1)/(t-1) ?

OpenStudy (anonymous):

\[\left(\begin{matrix}(t-1)(t^{2}+t+1) \\(t-1)\end{matrix}\right) = \left( t^{2}+t+1 \right)\]

OpenStudy (anonymous):

now u can do the integration

OpenStudy (anonymous):

ho did u get (t^{2}+t+1)?

OpenStudy (anonymous):

(a^3-b^3) = (a-b)*(a^2+ab+b^2)

OpenStudy (anonymous):

Best way to solve this problems is to find the roots of the polinomial and then simplificate the equation. Sheg did it pretty well

OpenStudy (anonymous):

ok..i got the answer!!thankz..